OpenStudy (anonymous):

if a + b=2 and a^2+ b^2= 5 then what is the value of a^3 + b^3?

OpenStudy (anonymous):

i think you are going to use \(a^3+b^3)=(a+b)(a^2+ab+b^2)\) at the moment you know \((a+b)\) and \(a^2+b^2\) so all that is left to know is \(ab\) any idea how to find it?

OpenStudy (anonymous):

i made a mistake there, it is \[a^3+b^3=(a+b)(a^2-ab+b^2)\] but the idea is the same, you need \(ab\)

OpenStudy (anonymous):

if you want a hint, to find \(ab\) note that \((a+b)^2=a^2+b^2+2ab\) you know all of those numbers except \(2ab\)

OpenStudy (anonymous):

wait? can you explain again? do i solve for a and b using the other formulas?

OpenStudy (anonymous):

you know that \(a+b=4\), and that \(a^2+b^2=5\)

OpenStudy (anonymous):

and also that \(a^3+b^3=(a+b)(a^2-ab+b^2)\) so \(a^3+b^3=4\times (5-ab)\)

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

you cannot solve for \(a\) and for \(b\) but you do not need to. you need to solve for \(ab\)

OpenStudy (anonymous):

it is not 4, it is 2

OpenStudy (anonymous):

ok my fault. but no matter, you have \[a^3+b^3=2(5-ab)\]

OpenStudy (anonymous):

what you need to finish is \(ab\)

OpenStudy (anonymous):

since \((a+b)^2=a^2+b^2+ab\) you have \(2^2=5+ab\)

OpenStudy (anonymous):

so would it be 5=2(5-ab)?

OpenStudy (anonymous):

no the left hand side you do not know it is the \(a^3+b^3\) that you are looking for

OpenStudy (anonymous):

it is \[a^3+b^3=2(5-ab)\]

OpenStudy (anonymous):

never mind....yea i just saw what i was doing...

OpenStudy (anonymous):

so since \((a+b)^2=a^2+b^2+2ab\) you have \[2^2=5+2ab\] \[4=5+2ab\] \[2ab=-1\] and so \(ab=-\frac{1}{2}\)

OpenStudy (anonymous):

making \[a^3+b^3=2(5+\frac{1}{2})\]

OpenStudy (anonymous):

(a+b)^2=a^2+b^2+2ab?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I see where you got it from.

OpenStudy (anonymous):

it is usually written as \[(a+b)^2=a^2+2ab+b^2\] but i wrote it differently as \[(a+b)^2=a^2+b^2+2ab\] so i could replace the \(a^2+b^2\) by 5

OpenStudy (anonymous):

but wouldnt it be -ab?

OpenStudy (anonymous):

would what be \(-ab\) ?

OpenStudy (anonymous):

oh you need \(-ab\) for sure, so answer your question but first we find \(2ab=-1\) and so \(-ab=\frac{1}{2}\)

OpenStudy (anonymous):

when you said it is usually written as (a+b)2=a2+2ab+b2 but i wrote it differently as (a+b)2=a2+b2+2ab

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

it is not \(-ab\) \[(a+b)^2=(a+b)(a+b)=a^2+2ab+b^2\]

OpenStudy (anonymous):

but where in this process does a^3 + b^3 go to?

OpenStudy (anonymous):

i think i have confused you

OpenStudy (anonymous):

we start with factoring \[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

you want to know what that is and so far we know \(a+b=2\) and \(a^2+b^2=5\) so we have most of it, we are just missing one piece we have \[a^3+b^3=2\times (5-ab)\]

OpenStudy (anonymous):

the only part that we don't know is \(-ab\)

OpenStudy (anonymous):

to find \(-ab\) we look at \[(a+b)^2\] which is \(2^2=4\) and notice that \[(a+b)^2=4=a^2+b^2+2ab\] we also know \(a^2+b^2=5\) so the above equation becomes \[4=5+2ab\]

OpenStudy (anonymous):

i see what you are saying but the only thing is where you get the (a+B)^2

OpenStudy (anonymous):

oh it was a gimmick to find \(-ab\) that is all a trick in other words i thought it up you have to find \(ab\) somehow

OpenStudy (anonymous):

all you know is \(a^2+b^2=5\) and \(a+b=2\) so you have to use what you got to get what you want

OpenStudy (anonymous):

Okay! It makes so much more sense now! Thanks!!!

OpenStudy (anonymous):

yw